OpenStudy (anonymous):

what is equivalent to -x squared plus 5x plus 46=18

OpenStudy (shamil98):

-x^2 + 5x + 46 = 18 -x^2 + 5x + 28 = 0 use the quadratic formula.

OpenStudy (anonymous):

So is that the answer?

OpenStudy (mertsj):

What did the problem ask?

OpenStudy (anonymous):

what equation is equivalent -x squared +5x+46=18